OpenStudy (anonymous):

the integral of sec 2x * tan 2x dx?

OpenStudy (amistre64):

whats the derivative of sec?

OpenStudy (rogue):

\[\int\limits_{}^{} \sec (2x) \tan (2x)dx\]\[u = 2x, \frac {1}{2}du = dx\]\[\frac {1}{2} \int\limits_{}^{} \sec u \tan u du = \frac {1}{2} \sec u = \frac {1}{2} \sec 2x\]

myininaya (myininaya):

+C

OpenStudy (amistre64):

the 2x pops out a 2 that needs to disappear so id say 1/2 sec(2x) yep

OpenStudy (amistre64):

+C lol

OpenStudy (rogue):

Always the constant... Damn.

myininaya (myininaya):

yes i contributed something lol

OpenStudy (anonymous):

Rogue, good job!

OpenStudy (amistre64):

lets try to avoid the "d" in the future please :)

OpenStudy (rogue):

I'll try ;)

OpenStudy (anonymous):

i like plus K

OpenStudy (anonymous):

so the antideriative of sec * tanx is secx?

OpenStudy (rogue):

Yup =]

OpenStudy (rogue):

+ C

OpenStudy (anonymous):

yes, anyway, it's not in my cal book, thank you